# Revisiting Direct Speech-to-Text Translation with Speech LLMs: Better Scaling than CoT Prompting?

### v1

@@ 1 Introduction
The dominant trend for creating S2TT resources is to translate existing ASR corpora into target languages [7, 8]. Following this approach, we construct a pseudo-labeled dataset through a translation pipeline (Figure 1), and train models with both Direct and CoT while scaling the amount of data. Our results demonstrate that Direct exhibits more stable scaling than CoT.
@@ 2.1 Translation pseudo-labeling
We generate pseudo-labeled data by translating the transcriptions of an ASR dataset (Figure 1). We use the same translation LLM employed as backbone for the S2TT model. To filter out generation errors or low-quality translations, we first score the samples with a Quality Estimation (QE) system. We use BLASER 2.0 [10] for this task because it provides reference-free, sentence-level translation scores that correlate strongly with human judgments. We compute the scores by evaluating the similarity of the transcription and translation with blaser_2_0_qe [^1] https://huggingface.co/facebook/blaser-2.0-qe and remove samples below $3.75$, which was chosen after inspecting samples in multiple languages. Despite estimating the quality of the translation, BLASER 2.0 is based on SONAR [11] embeddings, which have a shared representation space for speech and text across languages. For this reason, we next use a Language Identification (LID) system to discard possible targets in the wrong language. We use GlotLID v3 [12] to compute the probabilities of each of the $L$ considered languages $\bm{p}=(p_{1},...,p_{L})$, and remove samples in which the expected target language $p_{e}<0.5$.
@@ 2.2 Speech Translation Model
To fully leverage ASR and T2TT performance in S2TT, CoT models use prompts derived from the template shown in Figure 2. The same template is applied during inference, and generation is performed with beam search using five beams.
@@ 3.1 Data
Our training framework includes four types of datasets: ASR, T2TT, S2TT and pseudo-labeled S2TT (S2TTpl) in six languages: Catalan (ca), German (de), English (en), Spanish (es), French (fr), and Italian (it). Table 1 reports the amount of data per language. For ASR, we use the train splits of Common Voice 21.0 [17] and Multilingual LibriSpeech [18], totaling approximately 6,000 and 48,900 hours, respectively. For T2TT, we use Wikimedia [19] samples with 5 to 100 words, filtered using the same pipeline applied to pseudo-labeled S2TT data. For S2TT, we consider the train splits of Europarl-ST v1.1 [20] and CoVoST 2 [21], totaling around 630 and 1,600 hours, respectively. Finally, S2TTpl data is generated by translating all Common Voice 21.0 samples into all possible target languages. We evaluate our models in CoVoST 2 and Fleurs [22] test sets.
Table 1: Amount of data per language and total target tokens (considering transcriptions and translations as target in S2TT). The number of hours for S2TTpl corresponds to repeated utterances (one for each of the five target languages).
@@
Languages as Source (hours)
@@
Languages as Target (samples)
@@ 3.2 Experiments
To analyze the impact of scaling S2TT data, we incrementally add S2TTpl samples to each baseline, training 5 new systems with 20%, 40%, 60%, 80% and 100% of the full dataset respectively. We refer to these models as DirectaugXX and CoTaugXX, where xx is the percentage of pseudo-labeled data used. These models are trained independently, they are not checkpoints of a single experiment.
In the CoT setting, each speech–transcription pair in the pseudo-labeled data is repeated six times, one for each target language. We hypothesize this repetition may bias the model and harm the ASR performance, which is particularly critical for CoT prompting. To validate this, we train two variants of CoTaugXX: in CoTaugXX, the transcription step contributes to the loss, whereas in CoTaugXX it does not.
@@
Fig. 3: Performance on Fleurs (xcomet) when scaling the amount of S2TTpl data. Results are averaged across all the considered language directions.
@@ 3.3 Training Details
Both stage 1 and stage 2 last one epoch and employ AdamW optimizer, a cosine learning rate scheduler, and gradient clipping with a maximum norm of $1.0$.
In stage 1, we use a learning rate of $7\cdot 10^{-5}$, a warm up during the first $3\%$ of updates, and set a maximum sequence length of 1024. Tokens from different sequences are packed into full-length sequences to improve efficiency, following [13]. Training is run on 16 GPUs with a per-device batch size of 64, yielding an effective batch size of 256.
@@ 4 Results
CoT degrades ASR as data scales  Figure 3(b) shows that WER remains stable with Directaug across all S2TTpl data scales, whereas both CoTaug and CoTaug gradually increase it by up to +7.4+7.4% and +13+13% relative to their baseline. These trends correlate with the steady decrease of S2TT performance, reflecting their critical reliance on the ASR sub-task. This is also observed at 2020% of S2TTpl data, where CoTaug reaches its minimum WER alongside its highest xcomet score. Additionally, CoTBase shows a WER 22% lower than its counterpart strategy. This confirms the findings of [8], indicating that direct S2TT data improve ASR performance.
@@
Fig. 4: S2TT results evaluated on Fleurs test set with xcomet. The x-axis represents the percentage of the S2TTpl data with which the different Direct and CoT models were trained.

### v2

@@ 1 Introduction
The dominant trend for creating S2TT resources is to translate existing ASR corpora into target languages [17, 22]. Following this approach, we construct a pseudo-labeled dataset using a translation pipeline (Figure 1), and train models with both Direct and CoT while scaling the amount of data. Our results demonstrate that Direct exhibits more stable scaling than CoT.
@@ 2.1 Translation pseudo-labeling
We generate pseudo-labeled data by translating the transcriptions of an ASR dataset (Figure 1). We use the same translation LLM employed as backbone for the S2TT model (see §2.2). To filter out generation errors or low-quality translations, we first score the samples with a Quality Estimation (QE) system. We use BLASER 2.0 [4] for this task because it provides reference-free, sentence-level translation scores that correlate strongly with human judgments. We compute the scores by evaluating the similarity of the transcription and translation with blaser_2_0_qe [^1] https://huggingface.co/facebook/blaser-2.0-qe and remove samples below $3.75$, which was determined through inspection of samples across multiple languages. Although BLASER 2.0 estimates translation quality, it is based on SONAR [7] embeddings, which encode speech and text from multiple languages into a shared space. Thus, it does not explicitly enforce the intended target language. We therefore apply a Language Identification (LID) system, using GlotLID v3 [15], to compute language probabilities $\bm{p}=(p_{1},...,p_{L})$ and discard samples with expected language probability $p_{e}<0.5$.
@@ 2.2 Speech Translation Model
We use the prompts defined in Figure 2 for S2TT. We derive the templates from CoT for ASR and T2TT, which allows better integration of model capabilities from these tasks into S2TT. The same template is applied during inference, and generation is performed with beam search using five beams.
@@ 3.1 Data
Our training framework includes four types of datasets: ASR, T2TT, S2TT and pseudo-labeled S2TT (S2TTpl) in six languages: Catalan (ca), German (de), English (en), Spanish (es), French (fr), and Italian (it). Table 1 reports the amount of data per language. For ASR, we use the train splits of Common Voice 21.0 [1] and Multilingual LibriSpeech [19], totaling approximately 6,000 and 48,900 hours, respectively. For T2TT, we use Wikimedia [21] samples with 5 to 100 words, filtered using the same pipeline applied to pseudo-labeled S2TT data. For S2TT, we consider the train splits of Europarl-ST v1.1 [14] and CoVoST 2 [23], totaling about 630 and 1,600 hours, respectively. Finally, S2TTpl data is generated by translating all Common Voice 21.0 samples into all possible target languages. We evaluate our models in CoVoST 2 and Fleurs [3] test sets.
Table 1: Amount of data per language and total target tokens (considering transcriptions and translations as target in S2TT). The number of S2TTpl hours reflects repeated utterances, as each utterance is paired with five target languages.
@@
Hours per source language
@@
Samples per target language
@@ 3.2 Experiments
To analyze the impact of scaling S2TT data, we incrementally add S2TTpl samples to each baseline, training 5 new systems with 20%, 40%, 60%, 80% and 100% of the pseudo-labeled dataset. We refer to these models as DirectaugXX and CoTaugXX, where xx is the percentage of pseudo-labeled data used. These models are trained independently, they are not checkpoints of a single experiment.
In the CoT setting, each speech–transcription pair in the pseudo-labeled data is repeated six times, one for each target language. We hypothesize this repetition may harm the ASR performance, which is particularly critical for CoT prompting. To validate this, we train two variants of CoTaugXX: in CoTaugXX, the transcription step contributes to the loss, whereas in CoTaugXX it does not.
@@
Fig. 3: Performance on the Fleurs benchmark when scaling the amount of S2TTpl data. S2TT and T2TT are evaluated with xcomet ($\uparrow$) and ASR with WER ($\downarrow$). Note that $0\%$ corresponds to base variants and the rest to augXX. Results are averaged across all the considered language directions.
@@ 3.3 Training Details
Both stage 1 and stage 2 (see §2.2) last one epoch and employ AdamW optimizer, a cosine learning rate scheduler, and gradient clipping with a maximum norm of $1.0$.
In stage 1, we use a learning rate of $7\cdot 10^{-5}$, a warm up during the first $3\%$ of updates, and set a maximum sequence length of 1024. Tokens from different sequences are packed into full-length sequences to improve efficiency, following [25]. Training is run on 16 GPUs with a per-device batch size of 16, yielding an effective global batch size of 256.
@@ 4 Results
CoT degrades ASR as data scales  Figure 3(b) shows that WER remains stable with Directaug across all S2TTpl data scales, whereas both CoTaug and CoTaug gradually increase it by up to +7.4+7.4% and +13+13% relative to their baseline. These trends correlate with the steady decrease of S2TT performance, reflecting their critical reliance on the ASR sub-task. This is also observed at 2020% of S2TTpl data, where CoTaug reaches its minimum WER alongside its highest xcomet score. Additionally, CoTBase shows a 2-point WER reduction compared to its counterpart strategy. This confirms the findings of [22], indicating that direct S2TT data improve ASR performance.
@@
Fig. 4: S2TT results evaluated on Fleurs test set with xcomet. The x-axis represents the percentage of the S2TTpl data with which the different Direct and CoT models were trained. Note that $0\%$ corresponds to base variants and the rest to augXX.